the length and the width of a rectangle are in the ratio 3:2. if the length is increased by 15 and the width is decreased by 5, the ratio of the length to the width becomes 3:1. find the length and the width
Answer by AAfter.Com(18)

Let the length and width of the rectangle be l and w units
Given,
l/w = 3/2 => l = 3w/2
and (l + 15)/(w - 5) = 3/1 => l + 15 = 3w - 15
=> 3w = l + 30 = 3w/2 + 30
=> 3w/2 = 30
=> w = 20
=> l = (3x20)/2 = 30
Hence, the length and breadth of the rectangle are 30 and 20 units